OpenStudy (anonymous):

What will the last payment be to pay off a $1,250 computer plus 8% sales tax with monthly payments of $120?   $10 $11 $12 $20 $30

OpenStudy (anonymous):

First, we should figure out how much this computer actually costs. If you add an 8% sales tax, then you are adding 8% of the 1250 to the original 1250. So, 1250 + 8%(1250) = 1250 + (.08)(1250) = 1382.40. If you pay $120 per month times x months, then if you set that equal to 1382.40, the x will equal how many months it would take to pay for it. 120x = 1382.40 x = 11.52 However, we can only think in whole numbers (1, 2, 3, 4, etc.), not decimals. So we have to round. We can't round down because we won't pay for all the $1382.40 in 11 months. So, it'll take 12 months to pay it off.